given that\\(A=\\begin{pmatrix}1 & 2 & 3\\\\ 4 & 5 & 6 \\end{pmatrix}\\)\nand \\[B=\\begin{pmatrix} 1 & 2\\\\ 3& 4\\\\ 5& 6 \\end{pmatrix}\\]\n. Find AB
1
Expert's answer
2019-09-23T10:57:46-0400

"\begin{pmatrix}
 1 & 2 & 3 \\
 4 & 5 & 6
\end{pmatrix}" * "\begin{pmatrix}
 1 & 2 \\
 3 & 4 \\
 5 & 6
\end{pmatrix}" = "\begin{pmatrix}
 1*1 + 2*3 + 3*5 & 1*2 + 2*4 + 3*6 \\
 4*1 + 5*3 + 6*5 & 4*2 + 5*4 + 6*6
\end{pmatrix}" = "\begin{pmatrix}
 22 & 28 \\
 49 & 64
\end{pmatrix}"